Clare has two savings accounts. One account has a 5% interest rate and the other has a 6% rate. The amount invested in the 5% account is $1,000 more than the amount unvested in the 6% account. Together these accounts pay her $490 in interest yearly. How much does she have invested in each account?
Answer by ptaylor(1894)

Interest(I)=Principal(P)* Rate(R)* Time(T) or I=PRT
Let x=amount invested in the 6% account
Interest at 6%=x*0.06*1 or 0.06x
Then x+1000=amount invested in the 5% account
Interest at 5%=(x+1000)*0.05*1 or 0.05(x+1000)
Now we are told that together these accounts pay her $490 per year. So:
0.06x+0.05(x+1000)=490 get rid of parens
0.06x+0.05x+50=490 subtract 50 from both sides
0.06x+0.05x+50-50=490-50 collect like terms
0.11x=440 divide both sides by 0.11
x=$4000-------------------------amount invested at 6%
x+1000=$4000+$1000=$5000--------------amount invested at 5%
CK
0.06*4000+0.05*5000=490
240+250=490
490=490
Hope this helps-----------------ptaylor